OpenStudy (anonymous):

Find any intercepts:

OpenStudy (anonymous):

y = \[\frac{ x - 1 }{ x - 2 }\]

hartnn (hartnn):

to get x intercept, plug in y=0 in that equation

hartnn (hartnn):

to get y intercept, plug in x=0 in that equation

hartnn (hartnn):

@amelapal what did u get after plugging in y=0 in that equation ?

OpenStudy (anonymous):

When I plugged 0 for x, I got y = 1/2 but I'm not sure if I cross multiply when I plug in 0 for y.

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

y= 1/2 is correct! :) now 0 = (x-1)/(x-2) multiply x-2 on both sides, what u get ?

OpenStudy (anonymous):

x -1 = 0?

hartnn (hartnn):

correct! that would make x=1 as your x-intercept (1,0) got it ?

OpenStudy (anonymous):

yeah x = 1! Yes thank you!!

hartnn (hartnn):

welcome ^_^